OpenStudy (anonymous):

If f(x) = |x 4| + 5 find f(1)

OpenStudy (pfenn1):

Is the equation \[f(x) =\left| x^4 \right|+5\]

OpenStudy (anonymous):

nope its if f(x) = 8 3|x + 2| find f(6).

OpenStudy (pfenn1):

Is there an operator between the 8 and the 3?

OpenStudy (anonymous):

yes 8-3

OpenStudy (pfenn1):

so the equation is\[f(x)=8-3\left| x+2 \right|\]? Where \[\left| x+2 \right|\]

OpenStudy (pfenn1):

*is the absolute value of x+2?

OpenStudy (pfenn1):

So to determine the solution for f(6), you plug 6 in the equation for x.

OpenStudy (anonymous):

if [ f(x) = 8 - 3|x + 2| find f(6).

OpenStudy (pfenn1):

\[f(6)= 8-3\left| 6+2 \right|=?\]